OpenStudy (beccab003):

Can someone please help me with this problem? I don't understand it and I would really appreciate the help. I will medal the person with the best answer. Thanks! Sally decides to do an experiment to see how many times her MP3 player will shuffle through her songs and play her favorite group, “The Googly Girls” per hour. Her results are shown in the probability distribution table below. Find the probability that the MP3 player will play 1 or 2 songs by “The Googly Girls” each hour. x 0 1 2 3 p(x) 0.13 0.67 0.14 0.06 0.14 0.19 0.67 0.81

OpenStudy (amoodarya):

p(x=1 or 2)=p(x=1)+p(x=2)-p(x=1 and x=2) so p(x=1 and x=2)=0 so p(x=1 or 2)=p(x=1)+p(x=2)-0 =0.67 +0.14

OpenStudy (beccab003):

So the answer is 0.81?

OpenStudy (beccab003):

Thanks amoodarya! Can you help me with one more similar one?

OpenStudy (amoodarya):

sure

OpenStudy (beccab003):

Thanks! Look at the attachment please!

OpenStudy (amoodarya):

e(x)=p(x1)*x1 +p(x2)*x2 +...

OpenStudy (amoodarya):

E(x)=1(1/36)+2(1/18)+3(1/18)+4(1/12)+5(1/9)+7(0)+….

OpenStudy (beccab003):

So I multiply all of them together? I got 43/36.

OpenStudy (beccab003):

Are you still there @amoodarya?

OpenStudy (amoodarya):

it is greater than 43/36 please show me your work

OpenStudy (beccab003):

Okay. All I did is multiply and add the numbers you gave me: E(x)=1(1/36)+2(1/18)+3(1/18)+4(1/12)+5(1/9)+7(0)+….

OpenStudy (beccab003):

What that what I was supposed to do?

OpenStudy (amoodarya):

E(x)=1/36+2(1/18)+3(1/18)+4(1/12)+5(1/18)+8/18+9/36+10/18+12/9+15/18+16/36+18/18+20/18+24/18+25/18+30/18+36/36

OpenStudy (beccab003):

Okay. I did it and came up with 11.65. I guess I did it wrong again.

OpenStudy (beccab003):

I did it again and came up with 12.277777778

OpenStudy (amoodarya):

1/(36) [1+2+6+12+...) I think it 12

OpenStudy (beccab003):

It is 12? okay I guess I did it wrong. Are you pretty sure that that is it?

OpenStudy (amoodarya):

almost sure

OpenStudy (beccab003):

Okay thanks so much. I am taking a geometry course and in the very last section they started adding a bunch of statistics. Not nice. I was really confused by these two problems. Thanks so much for your assistance!

OpenStudy (amoodarya):

glad 2 help u

OpenStudy (beccab003):

:) I might need a little bit more help later today. Would it be alright if your on openstudy if i message or tag you? I don't want to do that if it will bug you though.